The length L of a triangle is twice as long as its with W and the area of the rectangle is 242 sq ft. how long is the rectangle?
Answer by mwaqas(2)


A=LW
W=L/2
A=L.L/2
242*2=L^2
484=L^2
TAking root both side
L=22Ft